OpenStudy (1018):

Express (3a + b + 3c, − a + 6b − c, 2a + b + 2c) as a linear combination of (3, − 1, 2) and (1, 6, 1).

OpenStudy (anonymous):

\[\langle 3a+b+3c,~-a+6b-c,~2a+b+2c\rangle=m\langle3,-1,2\rangle+n\langle1,6,1\rangle\] \[\begin{cases} 3a+b+3c=3m+n\\ -a+6b-c=-m+6n\\ 2a+b+2c=2m+n \end{cases}\] You want to solve for \(m\) and \(n\). Might I suggest a matrix method? I assume this is linear algebra, so you should be familiar with this. \[\begin{pmatrix} 3&1&3\\-1&6&-1\\2&1&2 \end{pmatrix}\begin{pmatrix}a\\b\\c\end{pmatrix}=\begin{pmatrix}3m+n\\-m+6n\\2m+n\end{pmatrix}\] or in augmented matrix form, \[\begin{pmatrix} 3&1&3&|&3m+n\\-1&6&-1&|&-m+6n\\2&1&2&|&2m+n\end{pmatrix}\] Row reduce to echelon form, then you have \[\begin{pmatrix} 1&0&1&|&m\\0&1&0&|&n\\0&0&0&|&0\end{pmatrix}\] So you have \(m=a+c\) and \(n=b\).

OpenStudy (1018):

Hey thanks a lot! Can I ask another question? Not a problem, just a question. :)

OpenStudy (anonymous):

Please do!